3.Among employees of a certain firm, 70% know C/C++, 60% know Java, and 50% know both languages. What portion of programmers
a. does not know Java?
b. does not know Java and C/C++?
c. knows C/C++ but not Java?
d. knows Java but not C++?
e. If someone knows Java, what is the probability that he/she knows C/C++?
f. If someone knows C/C++, what is the probability that he/she knows Java?

Found 2 solutions by jrfrunner, ValerieDavis:
Answer by jrfrunner(365)

3.Among employees of a certain firm, 70% know C/C++, 60% know Java, and 50% know both languages. What portion of programmers
a. does not know Java?
=1-0.6=0.4
b. does not know Java and C/C++?
know only C/C++=.7-.5=.2, know only java=.6-.5=0.1, know both=0.5
knows java or C/C++=.2+.5+.1=0.8
neither=1-.8=0.2
c. knows C/C++ but not Java?
.7-.5=0.2
d. knows Java but not C++?
.6-.5=0.1
e. If someone knows Java, what is the probability that he/she knows C/C++?
P(C/java)=P(both)/P(java)=0.5/.6=5/6
f. If someone knows C/C++, what is the probability that he/she knows Java?
P(java/C)=P(both)/P(C)=0.5/.7=5/7

Answer by ValerieDavis(6)

I am also a Java Programmer from Germany. This question are very usefull for all students or person wants to start career as a java developer. I'm stuck with this assignment in my do my project college.